OpenStudy (anonymous):

CAN I PLZ PLZ GET HELP I NEED IT BADLY I WILL GIVE U A MEDAL AND I WILL FAN U. AND THEIR ARE NO MULTIPLE CHOICES What is the slope of the following line? y = -4x + 2

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

@michealyoung can u help me with another one plz

OpenStudy (anonymous):

yea sure

OpenStudy (anonymous):

u need help

OpenStudy (anonymous):

Thxs so much just give me a sec

OpenStudy (anonymous):

What is the slope of the line graphed below?

rvc (rvc):

Hey NO DIRECT ANSWERS

OpenStudy (anonymous):

ok wen u hav a y=mx+b ur m is the slope

OpenStudy (anonymous):

5 over 5

OpenStudy (anonymous):

flutter u rvc

OpenStudy (anonymous):

@rvc please stay out of it thxs i just need help

OpenStudy (anonymous):

and thxs to @michealyoung i have help

OpenStudy (anonymous):

np

OpenStudy (anonymous):

@jazzy288a The Code of Conduct of the site clearly says not to give or ask for direct answers only. @rvc is correct in what they have said. http://openstudy.com/code-of-conduct

OpenStudy (michele_laino):

hint: for general equation of a line: y=m*x+n the slope is the coefficient m

OpenStudy (anonymous):

@michealyoung What is the y-intercept of the following line? y = 2x - 4

OpenStudy (anonymous):

ok my bad

OpenStudy (anonymous):

2

OpenStudy (anonymous):

im sorry rvc

OpenStudy (anonymous):

would u guys back off of @michealyoung

OpenStudy (anonymous):

Does the graphed line have a positive, negative, or zero slope?

OpenStudy (anonymous):

2 is NOT the y intercept of that last equation.

rvc (rvc):

Thanks Obey the rules :)

OpenStudy (anonymous):

ik i read it wrong an jus realized it

OpenStudy (anonymous):

y-mx+b M = slope B = y-intercept (positive if +, negative if -) So with an equation of y=2x-4 the slope is 2 and the y-intercept is...?

OpenStudy (anonymous):

For the positive, negative, or zero: Look at what the line is doing. A positive slope would go up towards the right. A negative slope would go down towards the right. A zero slope would be horizontal. What is your line doing? Which do you think it is?

OpenStudy (anonymous):

Solve the equation. |2h – 3| = 1 {1, 2} {-1, 2} {-2, 2}

OpenStudy (anonymous):

@sleepyjess can i get your hep plz